HIM 3322-COST CONCEPTS- NAME Joseph Castellano SEMI VARIABLE COS CORE $1200 C- February Agnil Stay ,820 4,420 2.750 luly 4,420 4,470 4,250 1,900 $3,720 Use the High-Low Method to detemine the Total Fised and Total Variabile Cosots (fhom the repair espense above) for the year 2. Use Regressios Analysis to determine the Total Fised and Toal Variuable Costs (from the sepair expen above) for the year. You must solve this problem using a spreasheet program 兵1140 TVC-」/.uing u า06 ,233-0,4 3 aa 78. 6a35> ver ycar

Just need the answer to number 2. Thanks
Answer #1

Take the data into excel.To perform regression analysis:

Goto Data->Analysis->Regression

Select Repair expenses for Y range and Rooms sold for X range and click OK

The new Summary sheet provides coefficient of X variable=1.0157310520362 and Intercept= 2198.5435520362

Hence the equation becomes Y=2198.5435520362+ 1.0157310520362X

Now, the total variable cosy TVC = Total rooms sold*1.0157310520362=1.0157310520362*23360= $3727.47738

And the total fixed cost TFC = Total cost - TVC= 50110-23727.47738= $6382.52262